OpenStudy (anonymous):

Find the difference of (−4x2 − 7x3 − 2) − (8x3 + 3x − 7x2). Be sure to show all of your work in a neat and organized fashion. Use the ^ (shift 6) symbol to show an exponent. For example, x squared can be written as x^2.

OpenStudy (ivettef365):

so you have -4x^2 -7x^3 -2 -8x^3 -3x +7x^2 now combine like terms

OpenStudy (anonymous):

hm

OpenStudy (anonymous):

I kinda need help with that ^^" Sorry think u could explain the method for me to find my answer?

OpenStudy (ivettef365):

combine like terms so you can combine -7x^3 - 8x^3 = -15x^3 combine -4x^2 + 7x^2 = 3x^2 all other terms are different, so you have -15x^3 + 3x^2 -3x -2

OpenStudy (anonymous):

so how did u get that answer @ivetef365?

OpenStudy (ivettef365):

What part you don't understand ??

OpenStudy (ivettef365):

first I got out of the parenthesis, then I combined liked terms

OpenStudy (anonymous):

well how u combined it

OpenStudy (ivettef365):

just add the numbers, you have -7x^3 -8x^3, so -7 -8 = -15, so it is -15x^3 same with -4x^2 + 7x^2 that is the same as 7-4 = 3 so it is 3x^2, when adding numbers with variables, just add the numbers the variable stays the same

OpenStudy (anonymous):

oh okay i think i got it now

OpenStudy (anonymous):

could u give me an example ?

OpenStudy (ivettef365):

ok :)

OpenStudy (anonymous):

o: